# Exposing a Fatal Flaw in Sample-based Quantum Diagonalization Methods

### v1

@@ 4.2 More correlated systems: Iron-sulfur clusters
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 49 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
@@
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 33 and Figure S10 therein). In Ref. 33, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 33, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. We believe this discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation53 reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF48. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible.
5 Conclusion
@@
{suppinfo}
Effect of probability density function scaling (Eq. (4)), N2 ground-state energy error as a function of the number of quantum samples, and CI probability distribution plots for N2 and [2Fe-2S] systems.

### v2

@@ IV.3 More correlated systems: Iron-sulfur clusters
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 51 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
@@
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 33 and Figure S10 therein). In Ref. 33, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 33, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. We believe this discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation[55] reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF[50]. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible. Notably, our results contradict the claim that for [2Fe-2S], quantum sampling is more efficient than the classic HCI method[33].
IV.4 Generality and Sampling Considerations
So far, we have considered two specific systems and a class of scaled distributions for sampling. With these systems, we found a clear trade-off between sampling efficiency and the compactness of the resulting CI expansion. Strengthening the generality of our results, in the Supporting Information, we provide further examples of 29 additional chemical systems derived from the W4-17-MR benchmark set[56] with some additional hand-picked systems (see Figs. S5-S34). The conclusions drawn so far appear to be valid across a range of chemical systems, with the typical obstacle being poor sampling efficiency (similar to N2). Of interest are systems such as B2 and twisted ethene. These systems exhibit a behavior similar to the [2Fe-2S] system in the sense that there is an intermediate region where QSCI can provide more compact CI expansions than HCI initialized from the HF determinant, although this advantage is eventually lost when reaching more sizeable CI expansions, i.e., more accurate energies. A common feature of these systems is that the Hartree-Fock determinant is a poor initial guess for HCI as it has negligible weight in the ground-state CI expansion. Therefore, when HCI is initialized from the HF determinant, there is an intermediate region with the possibility of converging to excited states. However, as the threshold $\varepsilon$ is lowered, eventually, one connects to determinants important for the ground state, which leads to rapid convergence to the ground-state energy. These drawbacks can sometimes be avoided with good chemical intuition (recall [2Fe-2S]) or starting the HCI from determinants selected from a CIS calculation (twisted ethene).
In our work, we have generally assumed that the sampling from QSCI should occur from the ground-state distribution, in line with the original proposal of QSCI[30, 31]. Arguably, this is also the implicit assumption in the larger-scale SQD studies[33], which have employed optimization-free LUCJ with parameter initialization from CCSD $t_{2}$-amplitudes for their quantum experiments. However, the authors note that “quantum-classical optimization could further improve the quality of the solutions.” and present in the supporting information simulator studies with LUCJ circuits “optimized to minimize the estimator energy” and perform in-depth analysis on the ground state wave function concentration for quantum sampling. Optimization (i.e., energy minimization) of such ansätze would, in all likelihood, bring the probability distributions closer to the FCI distribution. Our work has demonstrated clear sampling inefficiencies when employing such distributions. However, there are several ways in which the sampling inefficiency could be a non-issue. First, if future quantum devices turn out to be orders of magnitude faster than current hardware, brute-force sampling from ground-state distribution might be a viable strategy. Second, one could consider alternative sampling strategies. For a fixed number of determinants ($L$), a very efficient sampling strategy would likely involve sampling from a uniform distribution over the $L$ largest CI coefficients, requiring approximately $O(2L\log(L))$ samples to achieve near-full coverage[33]. While we cannot rule out the existence of accurate and sample-efficient quantum circuits, their practical design and implementation are unknown and have yet to be demonstrated for QSCI. It constitutes a highly non-trivial problem and would most likely require a solution to the FCI problem in the first place to find the $L$ largest coefficients. In contrast, sampling from the ground state distribution (or at least approximations thereof) seems feasible and has already been demonstrated to some extent.
V Conclusion
@@ VI Acknowledgement
VII Data availability statement
The data that support the findings of this study are available from the corresponding author upon reasonable request. Due to the large size of some datasets, specific arrangements may need to be made for data sharing. All details necessary for reproducing the results are provided in the manuscript and supplementary information. Scripts and input files for the CASCI, HCI, and QSCI calculations are available at https://github.com/peter-reinholdt/qsci-benchmarks.
VIII Supporting Information Available
Effect of probability density function scaling (Eq. (4)), N2 ground-state energy error as a function of the number of quantum samples, and CI probability distribution plots for N2 and [2Fe-2S] systems. Examples of QSCI performance for 29 additional chemical systems.